OpenStudy (anonymous):

Find the slope of every line that is parallel to the graph of the equation. y-x=3

OpenStudy (anonymous):

would the answer be 1?

OpenStudy (amistre64):

ax+by=c such that slope = -a/b

OpenStudy (amistre64):

the slope is 1

OpenStudy (akshay_budhkar):

yes.. y=mx+c y=x+3 where m is slope you can compare..

OpenStudy (anonymous):

it said i was wrong tho :(

OpenStudy (akshay_budhkar):

no dont worry.. the answer key is wrong. or you copied the question wrong

OpenStudy (amistre64):

then either the question is posted wrong; or the computer is programed to give the wrong answer

OpenStudy (anonymous):

well then how about the question x-3y=5 what would the slope be there?

OpenStudy (amistre64):

-a/b still; doesnt really change

OpenStudy (amistre64):

-1/-3 = 1/3

OpenStudy (anonymous):

it would be 1/3 right?

OpenStudy (akshay_budhkar):

Same way.. You will get the slope as my writing it is y=mx+C which gives 1/3